# On the discontinuities of Hausdorff dimension in generic dynamical Lagrange spectrum

### v1

@@ 3.6. Subhorseshoes and connection by periodic orbits
because $\Lambda^{n}\subset\Lambda_{t_{n}-\eta_{n}}$ and as $j_{r_{3}(n)}-j_{r_{2}(n)}>4N_{n}k-3N_{n}k=N_{n}k$ we have that $\sigma^{l}(x_{R})$ coincides with $\sigma^{l-j}(\theta_{n}^{1};\beta_{n}\theta_{n}^{2})$ in the central block of size $12N_{n}+$ 1 centered at the zero position.
@@
3.7. End of the proof of theorem 1.1 when the dimension is less than $1$
We are ready to obtain the desired contradiction. As the map $O$ takes only a finite number of different values, said $M$. Then by corollary 3.21 it would be impossible to have a sequence $n_{1}<n_{2}<...<n_{M+1}$ of elements of $\mathcal{I}$ such that for $i,j\in\{1,...,M+1\}$ with $i\neq j$$\Lambda^{n_{i}}$ and $\Lambda^{{n_{j}}}$ doesn’t connect before $\max\{t_{n_{i}},t_{n_{j}}\}$ in contradiction with proposition 3.14.
3.8. Proof of theorem 1.1 when the dimension is greater than or equal to $1$
Consider $\varphi\in\mathcal{U}^{*}$ such that $HD(\Lambda)\geq 1$, $f\in\mathcal{P}_{\varphi,f}$ and some closed sub interval $I\subset I_{\varphi,f}$ that doesn’t contain neither $c_{\varphi,f}$ nor $\tilde{c}_{\varphi,f}$. Observe that, in this case, by corollary 3.3, $\max L_{\varphi,f}=1$ and then for $t<\tilde{c}_{\varphi,f}$ one has $L_{\varphi,f}(t)<1$.
@@
As before, the set $P$ admits a decomposition $P=\bigcup\limits_{i\in\mathcal{I}}\tilde{\Lambda}_{i}$ where $\mathcal{I}$ is a finite index set and for any $i\in\mathcal{I}$$\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set. Note that if $i_{0},i_{1}\in\mathcal{I}$ are different and $\tilde{\Lambda}_{i_{0}}$ and $\tilde{\Lambda}_{i_{1}}$ are subhorseshoes, then $\tilde{\Lambda}_{i_{0}}$ and $\tilde{\Lambda}_{i_{1}}$ don’t connect before $\max I$.
@@
Observe that, as in the proposition 2.4, the first part of the theorem also holds for subhorseshoes of $\Lambda$ with Hausdorff dimension less than $1$. Therefore, if we set $c_{i}=\min\{x:x\ \textrm{is an accumulation point of}\ \ell_{\varphi,f}(\tilde{\Lambda}_{i})\}$, by proposition 3.7 there is some $i_{0}\in\mathcal{I}$ such that $c_{\varphi,f}=c_{i_{0}}$ and also by corollary 3.10 for any $i$ such that $c_{\varphi,f}<c_{i}$ the function $L_{i}$ doesn’t contribute with any discontinuity close $c_{i}$ to the discontinuity set of $L$ (note that it is possible to have $c_{i}\geq\max I$ for some $i$). Then, we conclude that $L_{\varphi,f}$ has finitely many discontinuities in the interval $I$ as we wanted to see.

### v2

@@ 3.6. Subhorseshoes and connection by periodic orbits
because $\Lambda^{n}\subset\Lambda_{t_{n}-\eta_{n}}$ and as $j_{r_{3}(n)}-j_{r_{2}(n)}>4N_{n}k-3N_{n}k=N_{n}k$ we have that $\sigma^{l}(x_{R})$ coincides with $\sigma^{l-j}(\theta_{n}^{1};\beta_{n}\theta_{n}^{2})$ in the central block of size $2N_{n}+1$ centered at the zero position.
@@
3.7. End of the proof of Theorem 1.1 when the dimension is less than 11
We are ready to obtain the desired contradiction. As the map $O$ takes only a finite number of different values, said $M$. Then by Corollary 3.21 it would be impossible to have a sequence $n_{1}<n_{2}<...<n_{M+1}$ of elements of $\mathcal{I}$ such that for $i,j\in\{1,...,M+1\}$ with $i\neq j$, $\Lambda^{n_{i}}$ and $\Lambda^{{n_{j}}}$ doesn’t connect before $\max\{t_{n_{i}},t_{n_{j}}\}$ in contradiction with Proposition 3.14.
3.8. Proof of Theorem 1.1 when the dimension is greater than or equal to 11
Consider $\varphi\in\mathcal{U}^{*}$ such that $HD(\Lambda)\geq 1$, $f\in\mathcal{P}_{\varphi,f}$ (see Proposition 3.2) and some closed subinterval $I\subset I_{\varphi,f}$ that doesn’t contain neither $c_{\varphi,f}$ nor $\tilde{c}_{\varphi,f}$. Observe that, in this case, by Corollary 3.3, $\max L_{\varphi,f}=1$ and then for $t<\tilde{c}_{\varphi,f}$ one has $L_{\varphi,f}(t)<1$.
@@
As before, the set $P$ admits a decomposition $P=\bigcup\limits_{i\in\mathcal{I}}\tilde{\Lambda}_{i}$ where $\mathcal{I}$ is a finite index set and for any $i\in\mathcal{I}$, $\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set. Note that if $i_{0},i_{1}\in\mathcal{I}$ are different and $\tilde{\Lambda}_{i_{0}}$ and $\tilde{\Lambda}_{i_{1}}$ are subhorseshoes, then $\tilde{\Lambda}_{i_{0}}$ and $\tilde{\Lambda}_{i_{1}}$ don’t connect before $\max I$.
@@
Observe that, as in Proposition 2.4, the first part of the theorem also holds for subhorseshoes of $\Lambda$ with Hausdorff dimension less than $1$. Therefore, if we set $c_{i}=\min\{x:x\ \textrm{is an accumulation point of}\ \ell_{\varphi,f}(\tilde{\Lambda}_{i})\}$, by Proposition 3.7 there is some $i_{0}\in\mathcal{I}$ such that $c_{\varphi,f}=c_{i_{0}}$ and also by Corollary 3.10 for any $i$ such that $c_{\varphi,f}<c_{i}$ the function $L_{i}$ doesn’t contribute with any discontinuity close to $c_{i}$ to the discontinuity set of $L_{\varphi,f}$ (note that it is possible to have $c_{i}\geq\max I$ for some $i$). Then, we conclude that $L_{\varphi,f}$ has finitely many discontinuities in the interval $I$ as we wanted to see.